OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#7779 — Optical loop: little crown
Scheduled Maintenance Report for Web Cloud
Completed
We have an optical fiber cut-off on the little crown loop , the sites IVR94 DEF92 MOL92 MAR92 BGN92 aren't secure anymore.
We identified the problem and opened a ticket with the fiber operator.

The service was not interrupted.

Update(s):

Date: 2012-12-18 10:16:14 UTC
Two welds were made in a box outside the main distribution frame. We lost 0.9dB and 1.3 dB. The signal remains in the allowed range. It is now secure.

Date: 2012-12-17 11:02:29 UTC
The suspected fault is a constraint on one of the strand.
We are waiting for FT to access in a closed local
in the NRA.

Date: 2012-12-17 10:12:51 UTC
An intervention is in progress.

The interruption between MIC92 and BGN92 on a single strand.
On BGN tests were performed to check the fiber.
From the unbundling room to RO, it is OK.

We will go on MIC92 to perform the same thing.
Posted Dec 17, 2012 - 00:27 UTC